will you help me solve this equation? -6b + 9 + 4b = -3b +14
Answer by stanbon(75887)

help me solve this equation?
-6b + 9 + 4b = -3b +14
Simplify each side:
-2b + 9 = -3b + 14
Add 3b to both sides:
b + 9 = 14
Subtract 9 from both sides:
b = 5
============
Cheers,
Stan H.
=======================